Expected Cash Flows


Year Project A Project B

0 -500 -875

1 100 150

2 110 200

3 120 252

4 175 375

5 240 530

6 300 680


a. If you were told that each project’s cost of capital was 12%, which project should be selected using the NPV criteria?

b. What is the profitability index for each project if the cost of capital is 12%?

c. What is the regular payback period for these two projects?


1
Expert's answer
2021-11-15T10:10:16-0500

a. If you were told that each project’s cost of capital was 12%, then:

"NPVa = -500 + 100\/1.12 + 110\/1.12^2 + 120\/1.12^3 + 175\/1.12^4 + 240\/1.12^5 + 300\/1.12^6 = 161.78."

"NPVb = -875 + 150\/1.12 + 200\/1.12^2 + 252\/1.12^3 + 375\/1.12^4 + 530\/1.12^5 + 680\/1.12^6 = 481.3."

So, projectB should be selected using the NPV criteria.

b. The profitability index for each project if the cost of capital is 12% is:

"PIa = 161.78\/500 = 0.324" or 32.4%,

"PIb = 481.3\/875 = 0.55" or 55%.

c. The regular payback period is 5 years for project A and 4,5 years for project B, so is higher for project A.